Question 3: Variance Analysis (20 marks in total)

You have introduced a standard cost accounting system for FOL. The following standard costs have been developed for the Cattle drench.

Direct materials (1 kilogram)

$ 20

Direct Labour (.25 hours)

$ 15

Overhead (DLH basis)

$ 3

$ 38

Production and cost information for May was:

Actual direct material purchased

1200

kilograms

Actual direct materials issued

900

kilograms

Actual output

800

kilograms

Actual cost of materials purchased

$ 25,500

Actual direct labur rate

$ 60

per hour

Actual direct labour hours

280

Actual overhad costs

$ 4,500

Required

  1. Calculate the following standard cost variances for May and provide general journal entries to record the cost flows for May.

(a) materials price

(b) materials usage

(c) direct labour rate

(d) direct labour efficiency

(e) total overhead
Answer #1

(a) Material price variance = (20-25500/1200)*1200 =1500 unfavorable

(b) Material usage variance = (1*800 - 900)*20 = 2000 unfavorable

(c) Direct labor rate variance = (15/0.25 - 60)*280 = 0

(d) Direct labor efficiency variance = (800*0.25 - 280)*15/0.25 = 4800 unfavorable

(e) Total overhead variance = 3*0.25*800 - 4500 = 3900 unfavorable